# The Absolute Anabelian Geometry of Virtual Curves Arising from Sections of Arithmetic Fundamental Groups of Configuration Spaces

### v1

@@ 5. Characterization of Geometric Fundamental Groups
Let $\mathcal{F}=(i,\text{\rm Id},\text{\rm Id}):\mathcal{C}^{\prime}=(X^{\prime},Y,f^{\prime},k,\bar{k},t,s)\to\mathcal{C}=(X,Y,f,k,\bar{k},t,s)$ be a decuspidalization as in Proposition 5.3; $\mathcal{D}:(X^{\prime\prime},Y^{\prime\prime},f^{\prime\prime},k^{\prime\prime},\bar{k}^{\prime\prime},t^{\prime\prime},s^{\prime\prime})$ a pointed virtual curve; $\mathcal{G}\coloneqq(h_{X},h_{Y},h_{k}):\mathcal{D}\to\mathcal{C}$ an étale covering. Then there exists a commutative diagram
${{\mathcal{D}^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathcal{F}^{\prime}}}$$\scriptstyle{{\mathcal{G}^{\prime}}}$${{\mathcal{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathcal{G}}}$${{\mathcal{C}^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\mathcal{F}}}$${{\mathcal{C}}}$
@@
${\Pi_{\mathcal{D}^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathcal{F}^{\prime}}}$$\scriptstyle{\pi_{\mathcal{G}^{\prime}}}$${\Pi_{\mathcal{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathcal{G}}}$${\Pi_{\mathcal{C}^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathcal{F}}}$${\Pi_{\mathcal{C}}}$
@@
${\Pi_{\mathcal{D}^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathcal{F}^{\prime}}}$$\scriptstyle{\pi_{\mathcal{G}^{\prime}}}$${\Pi_{\mathcal{D}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathcal{G}}}$${\Pi_{\mathcal{C}^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathcal{F}}}$${\Pi_{\mathcal{C}}}$
@@
Let $k$ be a mixed characteristic local field of residue characteristic $p$; $X$ a smooth hyperbolic curve over $k$; $s:G_{k}\rightarrow\Pi_{X}$ a section. Then, in the notation of § 2, the discussion entitled “Configuration Spaces”, and Definition 3.1(iii):
(i) $[\text{\rm pr}^{2/1}_{X},s]$ and $[\text{\rm pr}^{1}_{X},s]$ are pointed virtual curves.
@@
Assertion (i), (ii) follow immediately [cf. [18], Remark 2.1.1]. Assertion (iii) follows from assertion (ii) and Proposition 5.3(iii). The condition of assertion (iv) implies that $H$ contains the kernel of $\pi_{\mathcal{F}_{X}}:\Pi_{[\text{\rm pr}^{2/1}_{X},s]}\to\Pi_{[\text{\rm pr}^{1}_{X},s]}$. Therefore, by Proposition 5.7, there exists a decuspidalization $\mathcal{F}:\mathcal{D}^{\prime}\to\mathcal{D}$ such that $H\simeq\Pi_{\mathcal{D}^{\prime}}$, $J\simeq\Pi_{\mathcal{D}}$, and $\mathcal{F}_{X}|_{H}$ may be identified with $\pi_{\mathcal{F}}$, relative to the above isomorphisms. Thus, by Proposition 5.3(iii), one concludes that $(\mathcal{F}_{X}|_{H})^{\text{\rm ab-t}}$ is an isomoprhism. ∎
@@
Let $k$ be a mixed characteristic local field of residue characteristic $p$; $X$ a smooth hyperbolic curve over $k$; $s:G_{k}\rightarrow\Pi_{X}$ a section; $l\neq p$ a prime number. Then for any open normal subgroup $H\subset\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ of admissible type corresponding to an étale covering $\mathcal{D}$ of $[\text{\rm pr}^{2/1}_{X},s]$,
@@
By Proposition 5.8(iii) and Definition 5.9, $H$ is the inverse image of a normal open subgroup $J\subset\Pi_{[\text{\rm pr}^{1}_{X},s]}$ by the morphism $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}\rightarrow\Pi_{[\text{\rm pr}^{1}_{X},s]}$. By Lemma 5.11(i) below, $J$ corresponds to an étale covering $Y\rightarrow X$ such that $J\simeq\Pi_{[\text{\rm pr}^{1}_{J},s_{H}]}$, where
@@
denotes the first projection, and $s_{H}:G_{k_{H}}\rightarrow\Pi_{X\times_{k}k_{H}}$ the restriction of $s$ to $G_{k_{H}}$. By Proposition 5.3, $H\simeq\Pi_{[f,s_{H}]}$, where $f$ denotes the morphism naturally induced by the first projection
@@
Denote $[\text{\rm pr}^{1}_{J},s_{H}]$ by $\mathcal{D}_{J}$. Then by Proposition 5.8(iv),
@@
$\text{\rm Rank}_{p}((\Pi_{\mathcal{D}_{J}})^{\text{\rm ab-t}})-\text{\rm Rank}_{l}((\Pi_{\mathcal{D}_{J}})^{\text{\rm ab-t}})$
Moreover, by Lemma 5.11(iii) below,
@@
This completes the proof of Proposition 5.10. ∎
@@
$1\rightarrow\Pi_{Y\times_{k}\bar{k}}\rightarrow\Pi_{Y}\rightarrow G_{k}\rightarrow 1$
@@
which follows from the fact that both $X$ and $Y$ are the interiors of log smooth curves over $k$, and [[5], Theorem 2]. To show assertion (ii), one may assume without loss of generality that $Y$ admits a $k$-rational point. Then assertion (ii) follows from Theorem 4.20(ii)(iii). Assertion (iii) follows from [[13], Lemma 1.1.4(ii)] ∎
@@
By Proposition 5.10 and Definition 5.12, any subgroup of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ that is of field extension type contains $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$, as well as the commutator of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$. Therefore, $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$ is of geometric type. We claim that any subgroup $J$ of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ that is of geometric type is contained in $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$. Indeed, consider the image $I$ of $J$ in $G_{k}\simeq\Pi_{[\text{\rm pr}^{2/1}_{X},s]}/\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$. Since $I$ is a topologically finitely generated closed normal subgroup of $G_{k}$, by the elasticity of $G_{k}$ [cf. [15], Theorem 1.7(ii)], $I$ is either open or trivial. On the other hand, $I$ is contained in the intersection of the images in $G_{k}$ of the open subgroups $H\subset\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ of field extension type, hence is of infinite index in $G_{k}$ (since $G_{k}^{\text{\rm ab-t}}$ is infinite). Thus, $J\subset\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$, as desired. Therefore, $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$ may be characterized as the maximal subgroup of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ that is of geometric type. ∎
@@ 6. Automorphisms of 2-Configuration Spaces
In § 4 and § 5, we discussed various ways to characterize various objects of geometric origin. In this section, we first construct the function field of a smooth, geometrically connected curve of type $(0,3)$ over either a mixed-characteristic local field or a number field from the virtual fundamental group associated to a Galois section of the curve. Then, by generalizing the previous construction, we construct various scheme-theoretic objects, such as $\mathbf{F}$-graphs (where $\mathbf{F}$ is the category of fields, [cf. § 2, the discussion entitled “Graphs and Categories”]) whose vertices map to function fields of curves of type $(0,n)$, where $n\geq 3$, over either a mixed-characteristic local field or a number field, from certain virtual fundamental groups.
@@
Let $k$ be a field of characteristic zero such that there exists a prime number $l$ such that $k$ is $l$-cyclotomically full; $\bar{k}$ an algebraic closure of $k$; $X$ a smooth hyperbolic curve over $k$; $s:G_{k}\rightarrow\Pi_{X}$ a section; $\text{\rm diag}_{X}\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$ the conjugacy class of cuspidal inertia subgroups determined by the decuspidalization $[\text{\rm pr}^{2/1}_{X},s]\to[\text{\rm pr}^{1}_{X},s]$ of pointed virtual curves [cf. Definition 5.4(i), Proposition 5.8]; $H$ the normal closed subgroup of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ generated by $\text{\rm diag}_{X}$. Then the following exact sequences
@@
It follows from Proposition 5.3(ii) that the first and the second exact sequences are isomorphic. It follows from Lemma 5.11 that the second and the third exact sequences are isomorphic, and that there exists a natural bijection between $\text{\rm Cusp}([\text{\rm pr}^{1}_{X},s])$ and the set of conjugacy classes of inertia subgroups of cusps of $X$. ∎
@@
Generally speaking, in the situation of Proposition 6.1, one cannot group-theoretically characterize the element $\text{\rm diag}_{X}\in\mathcal{C}$. However, in the special case of type $(0,3)$, one may show that every class of cuspidal inertia subgroups that is fixed by the conjugation action of $G_{k}$ may be regarded as the class that represents the diagonal.
@@
[…]
@@
(vi) Note that for any $f\in I$, the $h\in H$ that satisfies the commutativity property in (v) is unique. In particular, we obtain a homomorphism $u:I\to H$, whose kernel $J$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. If $s:G_{k}\to\Pi_{X}$ is a section, then elements of $J$ may be regarded as automorphisms of $[\text{\rm pr}^{2/1}_{X},s]$.
@@
The assertions of Lemma 6.4 follow immediately from various definitions involved. ∎
@@
Let $k$ be a field of characteristic zero; $X$ a smooth curve over $k$ that is isomorphic to $\mathbf{P}^{1}_{k}\setminus\{0,1,\infty\}$. Denote by $\mathbb{S}_{5,X}$ (respectively, $\mathbb{S}_{3,X}$; $\mathbb{S}_{4,X}$; $\mathbb{V}_{X}$) the group $G$ (respectively, $H$, $I$, $J$) constructed in Lemma 6.4. Denote by $F_{X}:\mathbb{S}_{4,X}\to\mathbb{S}_{3,X}$ the morphism $u$ constructed in Lemma 6.4.
@@
By Lemma 6.6, there exists a $k$-isomorphism $h:X\xrightarrow{\sim}Y$ if and only if there exists a $k^{\prime}$-isomorphism $f^{\prime}:X\times_{k}k^{\prime}\xrightarrow{\sim}Y\times_{k}k^{\prime}$ such that for any element $\alpha\in\text{\rm Gal}(k^{\prime}/k)$, $(Y\times_{k}\alpha)\circ f^{\prime}=f^{\prime}\circ(X\times_{k}\alpha)$. On the other hand, this condition is equivalent to the condition that for any element $\alpha\in\text{\rm Gal}(k^{\prime}/k)$,
@@
In the situation of Proposition 6.1, assume further that $k$ is a number field or a mixed-characteristic local field. Then the following hold:
@@
(v) Suppose that $X$ satisfies the equivalent conditions in (iv). Then elements in $\mathbb{S}_{4,X}$ induce natural automorphisms of the virtual variety $[\text{\rm pr}^{2/1}_{X},s]$, hence (by (iii)) determine a morphism
@@
Assertions (i), (ii), and (iii) follow immediately from Proposition 4.18(i), Theorem 4.20(ii), Theorem 5.1, Proposition 5.2, and Theorem 5.13. To show assertion (iv), note that $\text{\rm diag}_{X}\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$ is fixed by the conjugation action of $G_{k}$. Therefore, by Proposition 5.8(ii) and Proposition 6.1, it suffices to show that $X$ is isomorphic to $\mathbb{P}^{1}_{k}\setminus\{0,1,\infty\}$ if and only if the action of $G_{k}$ on $\text{\rm Cusp}([\text{\rm pr}^{1}_{X},s])$ by conjugation is trivial. By Proposition 6.1, it suffices to show that $X$ is isomorphic to $\mathbb{P}^{1}_{k}\setminus\{0,1,\infty\}$ if and only if the action of $G_{k}$ on the $\text{\rm Cusp}(X)$ by conjugation is trivial. This follows from Lemma 6.4(iv) and Corollary 6.7.
Assertion (v) follows immediately from Theorem 4.20(ii), (iii); Lemma 6.4(v), (vi).
Next, we consider assertion (vi). Write $\{a_{1},a_{2},a_{3},a_{4}\}$ for the four elements of $\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$. For each $i\in\{1,2,3,4\}$, write $H_{i}$ for the subgroup of $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$ generated by the cuspidal inertia subgroups in $a_{i}$ (so $H_{i}$ is normal in $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$); $\Pi_{i}$ for the quotient $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}/H_{i}$; $\Delta_{i}$ for the kernel of the natural surjection $\Pi_{i}\twoheadrightarrow G_{k}$. For distinct $i,j\in\{1,2,3,4\}$, denote the image of $H_{j}$ in $\Pi_{i}$ by $H^{i}_{j}$. We may assume without loss of generality that $a_{1}=\text{\rm diag}_{X}$. Therefore, by Proposition 6.1, $\Pi_{1}$ is isomorphic to $\Pi_{X}$. We claim the following:
@@
First, we show existence. By (v), one may take $\theta_{j}^{i}$ to be the isomorphism induced by the natural action of $\sigma$ (considered as an element of $\mathbb{V}_{X}$ [cf. Lemma 6.4(vi)]) on $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$.
Uniqueness follows from the following claim, which follows immediately from Proposition 6.2:
[…]
@@
Now we have already reconstructed groups $\Pi_{i}$ and morphisms $\theta_{j}^{i}$ up to the composition with an inner automorphism. By Proposition 6.2, this data allows one to reconstruct four function fields that are isomorphic to $K(X)$ and isomorphisms between them. Finally, it follows from the condition that $\theta_{j}^{i}$ maps $H^{i}_{m}$ onto $H^{j}_{\sigma(m)}$ that these isomorphisms satisfy the desired composition relations. ∎
Finally, we consider the situation discussed in Proposition 6.8 in the case where $g=0$ and the action of the Galois group of the base field on the set of cusps is trivial.
@@
(ii) $(i,\text{\rm Id},\text{\rm Id}):[\text{\rm pr}^{1}_{X},s]\to[u,s]$ and $(h,\text{\rm Id},\text{\rm Id}):[\text{\rm pr}^{2/1}_{X},s]\to[v,s]$ are degree $1$ decuspidalizations [cf. Definition 5.4(ii)].
@@
$1\to\Delta_{[\text{\rm pr}^{2/1}_{Y},s^{\prime}]}\to\Pi_{[\text{\rm pr}^{2/1}_{Y},s^{\prime}]}\to G_{k}\to 1$
@@
Assertions (i) and (ii) follow immediately from the various definitions involved. Assertion (iii) follows immediately from Proposition 3.7.
@@
In the situation of Lemma 6.10, write $\alpha\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$ for the conjugacy class determined by $(h,\text{\rm Id},\text{\rm Id})$. Then the exact sequences

### v2

@@ 5. Characterization of Geometric Fundamental Groups
Let $\mathcal{F}=(i,\text{\rm Id},\text{\rm Id}):\mathcal{C}^{\prime}=(X^{\prime},Y,f^{\prime},k,\bar{k},t,s)\to\mathcal{C}=(X,Y,f,k,\bar{k},t,s)$ be a decuspidalization as in Proposition˜5.3; $\mathcal{D}:(X^{\prime\prime},Y^{\prime\prime},f^{\prime\prime},k^{\prime\prime},\bar{k}^{\prime\prime},t^{\prime\prime},s^{\prime\prime})$ a pointed virtual curve; $\mathcal{G}\coloneqq(h_{X},h_{Y},h_{k}):\mathcal{D}\to\mathcal{C}$ an étale covering. Then there exists a commutative diagram
[…]
@@
[…]
@@
[…]
@@
Let $k$ be a mixed characteristic local field of residue characteristic $p$; $X$ a smooth hyperbolic curve over $k$; $s:G_{k}\rightarrow\Pi_{X}$ a section. Then, in the notation of §˜2, the discussion entitled “Configuration Spaces”, and Definition˜3.1(iii):
(i) $\text{\rm pr}^{2/1}_{X},s]$ and $\text{\rm pr}^{1}_{X},s]$ are pointed virtual curves.
@@
Assertion (i), (ii) follow immediately [cf. [17], Remark 2.1.1]. Assertion (iii) follows from assertion (ii) and Proposition˜5.3(iii). The condition of assertion (iv) implies that $H$ contains the kernel of $\pi_{\mathcal{F}_{X}}:\Pi_{[\text{\rm pr}^{2/1}_{X},s]}\to\Pi_{[\text{\rm pr}^{1}_{X},s]}$. Therefore, by Proposition˜5.7, there exists a decuspidalization $\mathcal{F}:\mathcal{D}^{\prime}\to\mathcal{D}$ such that $H\simeq\Pi_{\mathcal{D}^{\prime}}$, $J\simeq\Pi_{\mathcal{D}}$, and $\mathcal{F}_{X}|_{H}$ may be identified with $\pi_{\mathcal{F}}$, relative to the above isomorphisms. Thus, by Proposition˜5.3(iii), one concludes that $(\mathcal{F}_{X}|_{H})^{\text{\rm ab-t}}$ is an isomoprhism. ∎
@@
Let $k$ be a mixed characteristic local field of residue characteristic $p$; $X$ a smooth hyperbolic curve over $k$; $s:G_{k}\rightarrow\Pi_{X}$ a section; $l\neq p$ a prime number. Then for any open normal subgroup $H\subset\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ of admissible type corresponding to an étale covering $\mathcal{D}$ of $\text{\rm pr}^{2/1}_{X},s]$,
@@
By Proposition˜5.8(iii) and Definition˜5.9, $H$ is the inverse image of a normal open subgroup $J\subset\Pi_{[\text{\rm pr}^{1}_{X},s]}$ by the morphism $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}\rightarrow\Pi_{[\text{\rm pr}^{1}_{X},s]}$. By Lemma˜5.11(i) below, $J$ corresponds to an étale covering $Y\rightarrow X$ such that $J\simeq\Pi_{[\text{\rm pr}^{1}_{J},s_{H}]}$, where
@@
denotes the first projection, and $s_{H}:G_{k_{H}}\rightarrow\Pi_{X\times_{k}k_{H}}$ the restriction of $s$ to $G_{k_{H}}$. By Proposition˜5.3, $H\simeq\Pi_{[f,s_{H}]}$, where $f$ denotes the morphism naturally induced by the first projection
@@
Denote $\text{\rm pr}^{1}_{J},s_{H}]$ by $\mathcal{D}_{J}$. Then by Proposition˜5.8(iv),
@@
$\text{\rm Rank}_{p}((\Pi_{\mathcal{D}_{J}})^{\text{\rm ab-t}})-\text{\rm Rank}_{l}((\Pi_{\mathcal{D}_{J}})^{\text{\rm ab-t}})$ .
Moreover, by Lemma˜5.11(iii) below,
@@
This completes the proof of Proposition˜5.10. ∎
@@
$1\rightarrow\Pi_{Y\times_{k}\bar{k}}\rightarrow\Pi_{Y}\rightarrow G_{k}\rightarrow 1.$
@@
which follows from the fact that both $X$ and $Y$ are the interiors of log smooth curves over $k$, and [[5], Theorem 2]. To show assertion (ii), one may assume without loss of generality that $Y$ admits a $k$-rational point. Then assertion (ii) follows from Theorem˜4.20(ii)(iii). Assertion (iii) follows from [[12], Lemma 1.1.4(ii)] ∎
@@
By Proposition˜5.10 and Definition˜5.12, any subgroup of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ that is of field extension type contains $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$, as well as the commutator of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$. Therefore, $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$ is of geometric type. We claim that any subgroup $J$ of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ that is of geometric type is contained in $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$. Indeed, consider the image $I$ of $J$ in $G_{k}\simeq\Pi_{[\text{\rm pr}^{2/1}_{X},s]}/\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$. Since $I$ is a topologically finitely generated closed normal subgroup of $G_{k}$, by the elasticity of $G_{k}$ [cf. [14], Theorem 1.7(ii)], $I$ is either open or trivial. On the other hand, $I$ is contained in the intersection of the images in $G_{k}$ of the open subgroups $H\subset\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ of field extension type, hence is of infinite index in $G_{k}$ (since $G_{k}^{\text{\rm ab-t}}$ is infinite). Thus, $J\subset\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$, as desired. Therefore, $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$ may be characterized as the maximal subgroup of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ that is of geometric type. ∎
@@ 6. Automorphisms of 2-Configuration Spaces
In §˜4 and §˜5, we discussed various ways to characterize various objects of geometric origin. In this section, we first construct the function field of a smooth, geometrically connected curve of type $(0,3)$ over either a mixed-characteristic local field or a number field from the virtual fundamental group associated to a Galois section of the curve. Then, by generalizing the previous construction, we construct various scheme-theoretic objects, such as $\mathbf{F}$-graphs (where $\mathbf{F}$ is the category of fields, [cf. §˜2, the discussion entitled “Graphs and Categories”]) whose vertices map to function fields of curves of type $(0,n)$, where $n\geq 3$, over either a mixed-characteristic local field or a number field, from certain virtual fundamental groups.
@@
Let $k$ be a field of characteristic zero such that there exists a prime number $l$ such that $k$ is $l$-cyclotomically full; $\bar{k}$ an algebraic closure of $k$; $X$ a smooth hyperbolic curve over $k$; $s:G_{k}\rightarrow\Pi_{X}$ a section; $\text{\rm diag}_{X}\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$ the conjugacy class of cuspidal inertia subgroups determined by the decuspidalization $\text{\rm pr}^{2/1}_{X},s]\to[\text{\rm pr}^{1}_{X},s]$ of pointed virtual curves [cf. Definition˜5.4(i), Proposition˜5.8]; $H$ the normal closed subgroup of $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ generated by $\text{\rm diag}_{X}$. Then the following exact sequences
@@
It follows from Proposition˜5.3(ii) that the first and the second exact sequences are isomorphic. It follows from Lemma˜5.11 that the second and the third exact sequences are isomorphic, and that there exists a natural bijection between $\text{\rm Cusp}([\text{\rm pr}^{1}_{X},s])$ and the set of conjugacy classes of inertia subgroups of cusps of $X$. ∎
@@
Generally speaking, in the situation of Proposition˜6.1, one cannot group-theoretically characterize the element $\text{\rm diag}_{X}\in\mathcal{C}$. However, in the special case of type $(0,3)$, one may show that every class of cuspidal inertia subgroups that is fixed by the conjugation action of $G_{k}$ may be regarded as the class that represents the diagonal.
@@
${X_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{\text{\rm pr}^{2/1}_{X}}$${X_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\text{\rm pr}^{2/1}_{X}}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$${X}$
@@
(vi) Note that for any $f\in I$, the $h\in H$ that satisfies the commutativity property in (v) is unique. In particular, we obtain a homomorphism $u:I\to H$, whose kernel $J$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. If $s:G_{k}\to\Pi_{X}$ is a section, then elements of $J$ may be regarded as automorphisms of $\text{\rm pr}^{2/1}_{X},s]$.
@@
The assertions of Lemma˜6.4 follow immediately from various definitions involved. ∎
@@
Let $k$ be a field of characteristic zero; $X$ a smooth curve over $k$ that is isomorphic to $\mathbf{P}^{1}_{k}\setminus\{0,1,\infty\}$. Denote by $\mathbb{S}_{5,X}$ (respectively, $\mathbb{S}_{3,X}$; $\mathbb{S}_{4,X}$; $\mathbb{V}_{X}$) the group $G$ (respectively, $H$, II, $J$) constructed in Lemma˜6.4. Denote by $F_{X}:\mathbb{S}_{4,X}\to\mathbb{S}_{3,X}$ the morphism $u$ constructed in Lemma˜6.4.
@@
By Lemma˜6.6, there exists a $k$-isomorphism $h:X\xrightarrow{\sim}Y$ if and only if there exists a $k^{\prime}$-isomorphism $f^{\prime}:X\times_{k}k^{\prime}\xrightarrow{\sim}Y\times_{k}k^{\prime}$ such that for any element $\alpha\in\text{\rm Gal}(k^{\prime}/k)$, $(Y\times_{k}\alpha)\circ f^{\prime}=f^{\prime}\circ(X\times_{k}\alpha)$. On the other hand, this condition is equivalent to the condition that for any element $\alpha\in\text{\rm Gal}(k^{\prime}/k)$,
@@
In the situation of Proposition˜6.1, assume further that $k$ is a number field or a mixed-characteristic local field. Then the following hold:
@@
(v) Suppose that $X$ satisfies the equivalent conditions in (iv). Then elements in $\mathbb{S}_{4,X}$ induce natural automorphisms of the virtual variety $\text{\rm pr}^{2/1}_{X},s]$, hence (by (iii)) determine a morphism
@@
Assertions (i), (ii), and (iii) follow immediately from Proposition˜4.18(i), Theorem˜4.20(ii), Theorem˜5.1, Proposition˜5.2, and Theorem˜5.13. To show assertion (iv), note that $\text{\rm diag}_{X}\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$ is fixed by the conjugation action of $G_{k}$. Therefore, by Proposition˜5.8(ii) and Proposition˜6.1, it suffices to show that $X$ is isomorphic to $\mathbb{P}^{1}_{k}\setminus\{0,1,\infty\}$ if and only if the action of $G_{k}$ on $\text{\rm Cusp}([\text{\rm pr}^{1}_{X},s])$ by conjugation is trivial. By Proposition˜6.1, it suffices to show that $X$ is isomorphic to $\mathbb{P}^{1}_{k}\setminus\{0,1,\infty\}$ if and only if the action of $G_{k}$ on the $\text{\rm Cusp}(X)$ by conjugation is trivial. This follows from Lemma˜6.4(iv) and Corollary˜6.7.
Assertion (v) follows immediately from Theorem˜4.20(ii), (iii); Lemma˜6.4(v), (vi).
Next, we consider assertion (vi). Write $\{a_{1},a_{2},a_{3},a_{4}\}$ for the four elements of $\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$. For each $i\in\{1,2,3,4\}$, write $H_{i}$ for the subgroup of $\Delta_{[\text{\rm pr}^{2/1}_{X},s]}$ generated by the cuspidal inertia subgroups in $a_{i}$ (so $H_{i}$ is normal in $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$); $\Pi_{i}$ for the quotient $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}/H_{i}$; $\Delta_{i}$ for the kernel of the natural surjection $\Pi_{i}\twoheadrightarrow G_{k}$. For distinct $i,j\in\{1,2,3,4\}$, denote the image of $H_{j}$ in $\Pi_{i}$ by $H^{i}_{j}$. We may assume without loss of generality that $a_{1}=\text{\rm diag}_{X}$. Therefore, by Proposition˜6.1, $\Pi_{1}$ is isomorphic to $\Pi_{X}$. We claim the following:
@@
First, we show existence. By (v), one may take $\theta_{j}^{i}$ to be the isomorphism induced by the natural action of $\sigma$ (considered as an element of $\mathbb{V}_{X}$ [cf. Lemma˜6.4(vi)]) on $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$.
Uniqueness follows from the following claim, which follows immediately from Proposition˜6.2:
($\star\star$) The image of the natural injection
@@
Now we have already reconstructed groups $\Pi_{i}$ and morphisms $\theta_{j}^{i}$ up to the composition with an inner automorphism. By Proposition˜6.2, this data allows one to reconstruct four function fields that are isomorphic to $K(X)$ and isomorphisms between them. Finally, it follows from the condition that $\theta_{j}^{i}$ maps $H^{i}_{m}$ onto $H^{j}_{\sigma(m)}$ that these isomorphisms satisfy the desired composition relations. ∎
Finally, we consider the situation discussed in Proposition˜6.8 in the case where $g=0$ and the action of the Galois group of the base field on the set of cusps is trivial.
@@
(ii) $(i,\text{\rm Id},\text{\rm Id}):[\text{\rm pr}^{1}_{X},s]\to[u,s]$ and $(h,\text{\rm Id},\text{\rm Id}):[\text{\rm pr}^{2/1}_{X},s]\to[v,s]$ are degree $1$ decuspidalizations [cf. Definition˜5.4(ii)].
@@
$1\to\Delta_{[\text{\rm pr}^{2/1}_{Y},s^{\prime}]}\to\Pi_{[\text{\rm pr}^{2/1}_{Y},s^{\prime}]}\to G_{k}\to 1.$
@@
Assertions (i) and (ii) follow immediately from the various definitions involved. Assertion (iii) follows immediately from Proposition˜3.7.
@@
In the situation of Lemma˜6.10, write $\alpha\in\text{\rm Cusp}([\text{\rm pr}^{2/1}_{X},s])$ for the conjugacy class determined by $(h,\text{\rm Id},\text{\rm Id})$. Then the exact sequences